OpenStudy (anonymous):

laplace inverse of 1/2(e^(-2x))((x^2-2x-1)/(x^3+x-2))

OpenStudy (anonymous):

The laplace transform or the inverse transform?

OpenStudy (anonymous):

the inverse laplace transform where t is greater than or equal to 2

OpenStudy (anonymous):

So they want you to find the transform and then calculate the transform that will get you back? Because that is not a function of t...?

OpenStudy (anonymous):

here is a print screen of the problem in the attachment. i honestly don't know what i'm doing! hope you can help!

OpenStudy (anonymous):

I can help! Find the laplace transform: The laplace transform is defined as: \[\int\limits_0^{\infty}f(t)e^{-st}dt\] Do you know how to evaluate integrals of this kind?

OpenStudy (anonymous):

do i separate the f(t) and e^(-nt)? and then take the limit of each as n goes to infinite?

OpenStudy (anonymous):

So this is what I would do, notice with t < 1 the function is zero, so your integral because: \[\int\limits_1^{\infty}(t^2-2t+2)e^{-st}dt\]

OpenStudy (anonymous):

becomes**

OpenStudy (anonymous):

but i thought at t=0 the function would equal 2...

OpenStudy (anonymous):

So evaluate each piece individually: \[\int\limits_1^{\infty}t^2e^{-st} - 2 \int\limits_1^{\infty} t e^{-st} dt +2 \int\limits_1^{\infty}e^{-st}dt\] Use the definition given.

OpenStudy (anonymous):

You can arbitrarily define a function to be anything at any given point, they are just declaring it to be that when t<1 /shrug

OpenStudy (anonymous):

Ready to work out this integral?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i thought the integral would be infinite...

OpenStudy (anonymous):

especially since e^x does not converge

OpenStudy (anonymous):

Its not e^x, its e^-st, the negative makes a huge difference. Okay, on the first integral use integration by parts: \[u=t^2, du=2t dt; dv = e^{-st}dt, v = -e^{-st}/s \implies \int\limits_1^{\infty}t^2e^{-st} \]\[= -t^2e^{-st}/2|_1^{\infty}+ 2 \int\limits_1^{\infty}te^{-st}dt\]

OpenStudy (anonymous):

Those should be over s actually.

OpenStudy (anonymous):

Perform integration by parts again and you're done. Let a be a finite value, evaluate the integrals and take the limits as a goes to infinity and you'll be done.

OpenStudy (anonymous):

so the function after solving the integral will be the answer to the inverse laplace?

OpenStudy (anonymous):

is s a variable or a constant. is t the only variable?